DURC 50m Open Rapid-Fire Competition Rules

1 Safety

All shooters are instructed to observe safe practise. Ensure your ammunition is secure before beginning the course of fire. Use an ammunition holder/block to do this. If you do not have one, one will be provided.

2 Targets

The target card will have three bulls aligned vertically as per the standard 100yard targets. The top card is for sighters, the bottom two for competition shots.

3 Course of Fire

There will be a five-minute preparation period during which competitors will get into position and DURC officals will setup/change target cards.

There will then follow a three-minute sighting period during which an unlimited number of sighting shots may be put in the top bull on the target card. At the end of this period, shooters should be ready to commence, with their ammunition arranged where it may be loaded rapidly and safely.

The Range Officer will ensure the firing line is ready, with all bolts/actions open and all breeches empty. The shooter may have the first round between his/her fingers if they so wish. The shooter MUST use an ammunition holder - rounds may not be laid on the mat as a safety precaution. During the detail, loading will be rapid and loose rounds will not be safe.

The Range Officer will then give the command “Commence”. Shooters will then have 90 seconds to put ten shots into the target card, five in the middle bull and five in the bottom bull. The Range Officer will give verbal warnings at 45 seconds elapsed, at twenty seconds to go, and a countdown from five seconds remaining. The Range Officer will then give the command “Stop, cease fire”. Any shooter who fires after this point is to be disqualified immediately. Should a shooter have a live round in the breech and the action/bolt closed, he should indicate so to the Range Officer who will observe the shooter fire the round into the backstop.

Shooters are not to break position until the Range Officer declares the line safe. The Range Officer will do so when all bolts/actions are opened and all breeched clear.

4 Competition arrangements

Ten shooters will shoot per detail. There will be no final. A tie for first, second or third place will be decided on by a shoot-off following the same course of fire. Any ties below third place will be broken by comparing first the number of tens in the shooter’s card, then the number of nines and so forth. Any remaining ties will not be broken.

5 Scoring

Scoring will be as per normal 50m cards, with the exception that a shot in the wrong bull on the shooters card will result in a deduction of one point from his overall score and he will lose any tie he/she places in by default unless the other shooters in that tie have also put a shot in the wrong bull.

Crossfires will result in the cross-firing shooter receiving the lowest score on the bull he cross-fired into. Where multiple cross-fires occour, the judgement of the Committee will decide the outcome.